1. You are considering a project that needs $50,000 investment to start. This is a 5 year project and you are expecting that the project generates annual cash flows after taxes of $10,000, $50,000, $90,000, $20,000 and -$30,000 respectively. What is the EAA of this project if the required rate of return is 8%? Round to the penny. Do not include the dollar sign. If there are multiple answers, then type NA.
Answer #1

To find EAA , firstly we need to find NPV

Using financial calculator,

Inputs: C0 = -50,000

C1 = 10,000. Frequency = 1

C2 = 50,000. Frequency = 1

C3 = 90,000. Frequency = 1

C4 = 20,000. Frequency = 1

C5 = -30,000. Frequency = 1

I = 8%

Npv = compute

We get, NPV = 67,854.20

Now, calculating EAA

EAA = (r × Npv) / (1-(1+r)^-n)

R = rate of interest= 8%

N = time = 5

Putting the values

= (0.08 × 67,854.2) / (1-(1+0.08)^-5)

= 5,428.34 / (1-(1.08)^-5)

= 5,428.34 / (1- 0.6806)

= 5,428.34 /0.3194

= 16,994.53